3.a. From the following uniformly distributed data scene calculate the probability w.r.t

statement.

The weekly output of a steel mill is a uniformly distributed random variable that lies

between 110 and 175 metric tons.

1. Compute the probability that the steel mill will produce more than 150 metric tons next

week.

2. Determine the probability that the steel mill will produce between 120 and 160 metric

tons "I WANT FULL LONG ANSWER"


1
Expert's answer
2021-09-10T06:21:01-0400

Question 1

P(150 <X< 175)

= (175-150)/(175-110)

= 25/65

= 0.385

Question 2

P(120 <X< 160)

= (160-120)/(175-110)

= 40/65

= 0.615